OpenStudy (anonymous):

Simplify \frac{ 12-3x^2 }{ x^2+5x-14 }=

OpenStudy (luigi0210):

Start off by factoring

OpenStudy (anonymous):

i got stuck

OpenStudy (anonymous):

i don't know how to factor the numerators

OpenStudy (anonymous):

the denominators are (x-2)*(x+7)

OpenStudy (anonymous):

is this correct ?

OpenStudy (luigi0210):

So far, now for the numerator, would it help if we did this: -(3x-12) Just factored out a negative

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

it's suppose to be 3x^2

OpenStudy (anonymous):

-3x^2

OpenStudy (luigi0210):

I took the negative out though..

OpenStudy (luigi0210):

And yea, sorry about that I forgot my exponent there Then factor out a 3.. -3(x^2-4)

OpenStudy (anonymous):

will this also work 3(4-x^2) ?

OpenStudy (luigi0210):

Yes, but you need to remove the negative sometime, because you won't be able to simplify anymore

OpenStudy (anonymous):

OK thanks

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

one more thing what do i do now ?

OpenStudy (luigi0210):

Do you know what (x^2-4) factors into?

OpenStudy (anonymous):

(x-2)^2

OpenStudy (anonymous):

ooo now i see it

OpenStudy (luigi0210):

Close, (x-2)^2=x^2-4x+4)

OpenStudy (luigi0210):

(x^2-4)=(x+2)(x-2)

OpenStudy (anonymous):

ok, i forgot,thanks

OpenStudy (luigi0210):

Anytime

OpenStudy (anonymous):

Luigi0210 i'm gonna ask more questions in the future,can i send you emails of those questions?

OpenStudy (luigi0210):

Just post them on here and tag me with the "@" and put my name

OpenStudy (luigi0210):

Like @andrijano_r.com

OpenStudy (anonymous):

like this @Luigi0210

OpenStudy (anonymous):

then the question

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Luigi0210 one more question Subtract and simplify \[\frac{ 9 }{ 2x^2+7x+5 }-\frac{ 3 }{ 2x^2+9x+10 }\] i calculated and got to here \[\frac{ 6x+15 }{ (2x+5)(x+1)(x+2) }\] is this correct? also help me i dont know what to do next .

OpenStudy (anonymous):

@Luigi0210